OpenStudy (anonymous):

2^x=4^y=8^z and xyz =288 then the value of1/2 x+1/4y+1/8z is

OpenStudy (anonymous):

2^x=2^2y=2^3z xyz=288

OpenStudy (anonymous):

2^(x+2y+3z)

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

\((2)^x = (2)^x\) \((4)^y = (2)^{2y}\) \((8)^z = (2)^{3z}\)

OpenStudy (anonymous):

\[(x)^a = (x)^b \implies a = b\]

OpenStudy (anonymous):

Meaning : \(x = 2y = 3z\)

OpenStudy (perl):

x = 2y x = 3z xyz = 288

OpenStudy (anonymous):

Still having problem anywhere?

OpenStudy (anonymous):

x*x/2*x/3=288 ,x^3=1728

OpenStudy (anonymous):

x=12

OpenStudy (anonymous):

so 1/24+1/96+1/288 =16/288 =1/18

OpenStudy (anonymous):

Got??

OpenStudy (anonymous):

Wait.. Have you done it correctly?

OpenStudy (anonymous):

nope.i think so

OpenStudy (anonymous):

\(x=12\), is right.. :)

OpenStudy (anonymous):

\[\frac{1}{2x} + \frac{1}{4y} + \frac{1}{8z} \implies \frac{1}{24} + \frac{1}{2(2y)} + \frac{1}{8 \frac{x}{3}}\]

OpenStudy (anonymous):

\(2y = x\), so \(2(x) = 24\)

OpenStudy (anonymous):

\(4y = 2(2y) = 2(x) = 24\) \[\frac{x}{3} = 4\]

OpenStudy (anonymous):

So, in denominators, all will be \(24\) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{1}{24} + \frac{1}{24} + \frac{1}{24} = ??\]

OpenStudy (anonymous):

3/24=1/8

OpenStudy (anonymous):

for x=3z 8z=?

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

\[x = 3z \implies z = \frac{x}{3}\]

OpenStudy (anonymous):

Getting?

OpenStudy (anonymous):

Now multiplying by \(8\) both the sides.. :)

OpenStudy (anonymous):

\[8z = 8 \times \frac{x}{3} = 8 \times \frac{\cancel{12}^4}{\cancel{3}} = 32\]

OpenStudy (anonymous):

Sorry, but you should understand that I am not God, I am man only.. :P

OpenStudy (anonymous):

lol:)

OpenStudy (anonymous):

\[\frac{1}{24} + \frac{1}{24} + \frac{1}{32} \]

OpenStudy (anonymous):

rofl::)

OpenStudy (anonymous):

Sorry, rofl?

OpenStudy (anonymous):

How much you got?

OpenStudy (anonymous):

11/96 is the right answer.:)

OpenStudy (perl):

x= 12 , y = 24, z= 36 1/(2*12 ) + 1 / ( 4 *24 ) + 1 / ( 8 * 36 ) = 1/18

OpenStudy (anonymous):

nope @perl.

OpenStudy (anonymous):

\[\frac{11}{96}\] I also got the same.. :)

OpenStudy (anonymous):

another two problems

OpenStudy (anonymous):

\[x = 12 = 2y \not \implies y = 24\]

OpenStudy (anonymous):

*does not imply.. Yeah sure, go ahead.. :)

OpenStudy (anonymous):

can i close the question or put here

OpenStudy (anonymous):

It is totally up to you.. :)

OpenStudy (anonymous):

simple one makes me time so .over here x+1/x=5 then x^6+1/x^6=?

OpenStudy (anonymous):

x+(1/x)=5 then x^6+(1/x^6)=?

OpenStudy (anonymous):

Okay wait.. :)

OpenStudy (anonymous):

Do you know the formula for : \((a+b)^3\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

a^3+b^3+3ab(a+b) .here i used it but it gave 3 as answer

OpenStudy (anonymous):

Let us go step by step, with the basic formula we have.. :)

OpenStudy (anonymous):

x^2+1/x^2=23

OpenStudy (anonymous):

\[(a+b)^2 = a^2 + b^2 + 2ab \implies \color{green}{a^2 + b^2 = (a+b)^2 - 2ab}\]

OpenStudy (anonymous):

Oh, you have used that?? Good.. :)

OpenStudy (anonymous):

\(x^2 + \frac{1}{x^2} = 25 - 2 = 23\)

OpenStudy (anonymous):

(x^2+1/x^2)^3

OpenStudy (anonymous):

23^3=12167

OpenStudy (anonymous):

\[(a+b)^3 = a^3 + b^3 + 3ab(a+b) \implies x^6 + \frac{1}{x^6} = (23)^3 - 3(23)\]

OpenStudy (anonymous):

12098

OpenStudy (anonymous):

\[\implies 23(23^2 - 3) = 23(529 - 3) = 23 \times 526 = ?\]

OpenStudy (anonymous):

\(12098\)

OpenStudy (anonymous):

Good.. Is that the right answer or not?

OpenStudy (anonymous):

another formulae also there to find see. (5)(a+b)^2

OpenStudy (anonymous):

Which formula?

OpenStudy (anonymous):

same (a+b)^3=(a+b)(a+b)^2

OpenStudy (anonymous):

Yes.. :)

OpenStudy (perl):

maybe I misread the question

OpenStudy (anonymous):

It is okay.. :)

OpenStudy (perl):

what was the correct answer ?

OpenStudy (anonymous):

And what is the \(\text{Third}\) question?

OpenStudy (anonymous):

5(23)^2-2 something wrong

OpenStudy (anonymous):

11/96 is correct @perl

OpenStudy (perl):

how do you get that? x = 12 , correct?

OpenStudy (anonymous):

I think you cannot use that formula here.. :)

OpenStudy (perl):

, x = 12, y = 24, z = 36

OpenStudy (anonymous):

x = 12 is correct.. :)

OpenStudy (anonymous):

'o' is the centre of the circle,AB is the chord of the circle OM perpendicular to AB.If AB=20cm and OM=2sqrt(11)cm,then the radius of the circle is,

OpenStudy (perl):

how did you get 11/ 96, i must be reading the formula incorrectly

OpenStudy (anonymous):

2y = x not 2x = y @perl

OpenStudy (anonymous):

you are finding y and z incorrectly.. :)

OpenStudy (perl):

ok

OpenStudy (perl):

y =6 and z = 3 then

OpenStudy (anonymous):

This looks good.. :)

OpenStudy (anonymous):

@harish1 while using that formula do mention what are a and b there.. :)

OpenStudy (anonymous):